I am reading a Turn and Talk question in a math book.
It says "How do you know when a product
will be less than one of its factors?"
I thought products are always equal to or greater than its factors.
If a product will be less than one of its factors, I need to see
an example.

Answer by greenestamps(13198)



If you are only working with natural numbers (counting numbers -- 1, 2, 3, ...) then the product of two numbers is always equal to or greater than each factor.

But if we allow fractions and one of the factors is less than 1, then the product of that number and a second factor will be less than the second factor. For example,

(1/2)*4 = 2, which is less than 4.

or

(1/2)*(2/3) = 1/3

There the product is less than either of the factors, because both factors are less than 1.

And there are other examples if we allow different kinds of numbers. For example, if the first factor is positive and the second is negative, then the product is negative; and a negative number is less than any positive number.

Example: (3)*(-2) = -6, which is less than 3